What is the buoyant force for a cube that displaces 0.530 kg of water when it is submerged in water?

What is the volume of the displaced water
Answer #1

Mass of water displaced = 0.530 kg

Since the buoyant force acting on the cube is given by Vdg

where V is volume of water displaced and d is the density of water.

Since Volume x density = mass of water displaced hence buoyant force = mg = 0.53*10 = 5.3N

now since volume x density = mass

and density of water = 1000kg/m^3

hence volume = 0.53/1000 = 5.3 x 10^-4 m^3